OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points)

OpenStudy (anonymous):

Something is missing please provide it.

OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 - 36 = 0 and x2 = 8x - 12 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

so for part 1 you have to solve x right?

OpenStudy (anonymous):

yea but idk how

OpenStudy (anonymous):

This ones easy, just take the 36 to the other side of the equals and take the square root. x^2 - 36 = 0 -> x^2 = 36 ----> x = +/- 6 (so it can be positive 6 or negative 6) Second question you need to put all the quadratic on one side and make it equal to zero. Then find two factors that add up to the middle term and multiply out to give the final term. One of the brackets must thus equal zero or both of them if the equation is true and these are your answers. x^2 = 8x - 12 ---> x^2 - 8x + 12 = 0 ---> (x-6)(x-2) = 0 so x = 6 and/or 2 Both of these solutions represent where the parabola intersects the x axis. Part 2: Draw the graphs and compare them. The x^-36 graph will be symmetric across the y axis whereas the other one will be symmetric across y = 4. Both graphs will be upwards pointing (e.g. a smile) Hope this helped! (:

OpenStudy (anonymous):

If that didnt help this is easier to understand: 2 – 36 = 0 and x2 = 8x – 12 x^2-36=0 difference of squares (x+6)(x-6) x^2-8x+12=0 (x-6)(x-2) it gets easier with experience

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Your welcome can I have a medal?